OpenStudy (anonymous):

What is the equation of a line that has a slope 4 and y intercept -3? A. Y=4x B. Y= 4x+32 C.y=4x - 3 D. 1/4x -3

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

y=mx+b where m is the slope and b is the y - intercept.

OpenStudy (anonymous):

4 - gradient & -3 ~ Y intercept so the formula is y = gradient * x + Y intercept y=mx + c

OpenStudy (anonymous):

therfore if u substitute the given numbers into the formula it is y=4x -3 and that is C! Hope it helped :)